OpenStudy (anonymous):

FIND THE INVERSE OF 3X+5/7

OpenStudy (anonymous):

i bet this is \[y=\frac{x+5}{7}\]

OpenStudy (anonymous):

izzit?

OpenStudy (anonymous):

NOPEEEE

OpenStudy (anonymous):

f−1(x) = The quantity of seven x minus five, over three f−1(x) = The quantity of seven x plus five, over three −f−1(x) = Seven over the quantity of three x plus five f−1(x) = Seven over the quantity of five minus three x ARE THE ANSWER CHOICES

OpenStudy (anonymous):

???

OpenStudy (anonymous):

\[f(x)=\frac{3x+5}{7}\] and you want \(f^{-1}(x)\) right?

OpenStudy (anonymous):

put \[x=\frac{y+5}{7}\] and solve the equation for \(y\) in two steps

OpenStudy (anonymous):

multiply by \(7\)to get \[7x=y+5\] then subtract \(5\) and get \[y=7x-5\] and that is your inverse